# Modeling Neural Activity with Conditionally Linear Dynamical Systems

### v1

@@ 1 Introduction
A central problem in neuroscience is to capture how neural dynamics are affected by external sensory stimuli, task variables, and behavioral covariates. To address this, a longstanding line of research has focused on characterizing neural dynamics through recurrent neural networks (RNNs) and their probabilistic counterparts, state-space models (SSMs; for reviews, see Paninski et al. 2010; Duncker & Sahani 2021; Durstewitz et al. 2023).
Early work in this area utilized latent linear dynamical systems (LDS) with Gaussian observation noise. Although these assumptions are restrictive, they are beneficial in two respects. First, they simplify probabilistic inference by enabling Kalman smoothing and expectation maximization (EM)—two classical and highly effective methods (see e.g., Ghahramani & Hinton 1996). Second, they produce models that are mathematically tractable to analyze with well-established tools from linear systems theory (Kailath, 1980). Indeed, many influential results in theoretical neuroscience have come from purely linear models (e.g. Seung 1996; Goldman 2009; Murphy & Miller 2009).
In reality, most neural circuits do not behave like time-invariant linear systems. Thus, more recent work from the machine learning community has cataloged a variety of nonlinear models for neural dynamics. Although these new models often predict held out neural data more accurately than LDS models, they are generally more difficult to fit and more difficult to understand. Thus, there has been a proliferation of competing architectures—e.g. RNNs (e.g. Pandarinath et al. 2018), transformers (e.g. Ye & Pandarinath 2021), and diffusion-based methods (Kapoor et al., 2024)—as well as competing training and probabilistic inference methods—e.g. generalized teacher forcing (Hess et al., 2023), amortized variational inference (e.g. Pandarinath et al. 2018), and sequential Monte Carlo (e.g. Pals et al. 2024). Choosing among these strategies and scientifically interpreting the outcomes is challenging.
@@
Finally, CLDS models have connections to several existing methods (§4). For example, they can be viewed as a dynamical extension of a Wishart process model (Nejatbakhsh et al., 2023) and an extension of Gaussian Process Factor Analysis (GPFA; Yu et al. 2009) with a learnable kernel and readout function that can vary across time and conditions. They are also similar to various forms of switching linear dynamical systems models (Petreska et al., 2011; Nassar et al., 2019; Hu et al., 2024). The key difference is that the “switching” in CLDS models is governed by an observed covariate vector, ${\bm{u}}_{t}$, rather than by a discrete latent process. This makes inference in the CLDS model much more straightforward, albeit at the price of not being fully unsupervised.
@@ 2 Methods
We use ${\mathbf{f}}(\cdot)\sim\mathcal{GP}^{N}(\bm{m}(\cdot),k(\cdot,\cdot))$ with mean $\bm{m}:\mathcal{X}\to\mathbb{R}^{N}$ and kernel $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, to denote samples ${\mathbf{f}}:\mathcal{X}\to\mathbb{R}^{N}$ obtained from stacking independent Gaussian processes into an $N$-dimensional vector.
@@ 2.1 Conditionally Linear Dynamical Systems
Consider an experiment with $N$ neurons recorded over $K$ trials of length $T$. Our dataset consists of the recorded neural trajectories $\{{\mathbf{y}}_{1:T}^{k}\}_{k=1}^{K}$, with ${\mathbf{y}}_{t}\in\mathbb{R}^{N}$ at time-step $t\in\{1,\dots,T\}$, along with the corresponding experimental conditions $\{{\bm{u}}_{1:T}^{k}\}_{k=1}^{K}$, with ${\bm{u}}_{t}$ in the condition space $\mathcal{U}$. By experimental conditions, we refer to available neural data covariates, either experimentally set or collected as measurements. These conditions, see Fig. 1a, can vary over time or remain constant (compare §3.3 vs. §3.4). They can also be a step function over time (e.g. animal moving vs. not moving), resulting in a switching-like mechanism between different dynamics.
We model response ${\mathbf{y}}_{t}$ as emissions from a latent time-varying linear dynamical system in ${\mathbf{x}}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{D}$, with the dynamics governed by the conditions ${\bm{u}}_{t}$. Specifically,
@@
evolving with time steps $t\in\{1,\dots,T\}$ from initial condition ${\mathbf{x}}_{1}\sim\mathcal{N}\left({\mathbf{x}}_{1};{\mathbf{m}}({\bm{u}}_{1}),Q_{1}\right)$, and where $\bm{\epsilon}_{t}\sim\mathcal{N}(0,Q)$ and $\bm{\omega}_{t}\sim\mathcal{N}(0,R)$ are sources of noise, both sampled i.i.d. over time. We assume that the latent variables ${\mathbf{x}}_{t}$ follow smooth dynamics defined by time-varying linear matrices ${\mathbf{A}}({\bm{u}})\in\mathbb{R}^{D\times D}$ from initial mean ${\mathbf{m}}({\bm{u}}_{1})\in\mathbb{R}^{D}$, with bias terms ${\mathbf{b}}({\bm{u}})\in\mathbb{R}^{D}$, ${\mathbf{d}}({\bm{u}})\in\mathbb{R}^{N}$, and emissions governed by ${\mathbf{C}}({\bm{u}})\in\mathbb{R}^{N\times D}$. See graphical depiction in Fig. 1b.
Conditions are typically treated as additive inputs, influencing the dynamics in (1a) via additive terms of the form ${\mathbf{B}}{\bm{u}}_{t}$ for a linear encoding matrix ${\mathbf{B}}\in\mathbb{R}^{D\times\left|\mathcal{U}\right|}$. Instead, the system in (1) parameterizes a family of linear systems indexed by a continuous time-varying variable, ${\bm{u}}_{t}$, which is observed. In fact, the system in (1) can be thought of as the linearization in ${\mathbf{x}}_{t}$ of a fully nonlinear system in ${\mathbf{x}}_{t}$ and ${\bm{u}}_{t}$, under additive noise—we explore this relationship in Appendix §A.2. The mapping of experimental conditions onto linear dynamics, ${\bm{u}}\mapsto\{{\mathbf{A}}({\bm{u}}),{\mathbf{b}}({\bm{u}}),{\mathbf{C}}({\bm{u}}),{\mathbf{d}}({\bm{u}}),{\mathbf{m}}({\bm{u}})\}$, is allowed to be nonlinear and learnable. Specifically, we place an approximate Gaussian Process (GP) prior on each entry of the parameters through a finite expansion of basis functions, leveraging regular Fourier feature approximations (Hensman et al., 2018). For any ${\mathbf{M}}\in\{{\mathbf{A}},{\mathbf{b}},{\mathbf{C}},{\mathbf{d}},{\mathbf{m}}\}$, we consider a prior over each $i,j$-th entry of the form:
@@
truncated at $L\in\mathbb{N}$ basis functions. Intuitively, each basis function $\phi_{\ell}:\mathcal{U}\to\mathbb{R}$ is fixed, and the randomness in the prior purely comes from the weights, ${\textnormal{w}}_{\ell}^{(ij)}$, which are drawn from a standard normal distribution. When constructed appropriately, the prior in equation (1c) converges to a nonparametric Gaussian process in the limit that $L\rightarrow\infty$. In our experiments, the basis functions $\{\phi_{\ell}\}_{\ell=1}^{L}$ are chosen and scaled as to approximate a GP prior of the form ${\mathbf{M}}_{ij}(\cdot)\sim\mathcal{GP}(0,k_{u})$ for the squared exponential kernel $k_{u}$ with variance $\sigma^{2}$ and length-scale $\kappa$ (Borovitskiy et al., 2020).
We denote ${\mathbf{F}}=\{{\mathbf{A}},{\mathbf{b}},{\mathbf{C}},{\mathbf{d}},{\mathbf{m}}\}$ as the set of random functions, and analogously the parameter set ${\mathbf{F}}({\bm{u}})=\{{\mathbf{A}}({\bm{u}}),{\mathbf{b}}({\bm{u}}),{\mathbf{C}}({\bm{u}}),{\mathbf{d}}({\bm{u}}),{\mathbf{m}}({\bm{u}})\}$ for any experimental condition ${\bm{u}}\in\mathcal{U}$. The model distribution
$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{1:T}\mid{\mathbf{A}},{\mathbf{b}},{\mathbf{C}},{\mathbf{m}},{\bm{u}}_{1:T})=\\ p({\mathbf{y}}_{1:T},{\mathbf{x}}_{1:T}\mid{\mathbf{F}},{\bm{u}}_{1:T})$ (2)
describes a time-varying latent Linear Dynamical System, conditioned on a parameter sequence set at experimental conditions. Therefore, we refer to the model in equations (1) as a Conditionally Linear Dynamical System (CLDS). Our CLDS implementation is available at:
https://github.com/neurostatslab/clds
@@ 2.2 CLDS modeling choices
A second source of flexibility is the encoding of experimental covariates, ${\bm{u}}$. Recall our notation from section 2.1, that ${\bm{u}}_{t}^{k}$ represents experimental covariates at time $t\in\{1,\dots,T\}$ and trial $k\in\{1,\dots,K\}$. A simple, and broadly applicable, modeling approach would be to set ${\bm{u}}_{t}^{k}=t$. This achieves a time-varying LDS model in which the GP prior encodes smoothness over time. This is similar in concept to fitting an linear model to data over a sliding time window (see e.g. Costa et al. 2019; Galgali et al. 2023). However, the CLDS formulation of this idea is fully probabilistic, which has several advantages. For example, we will see that one can use a single pass of Kalman smoothing to infer the distribution over the latent state trajectory, ${\mathbf{x}}_{1:T}^{k}$, within each trial. It is comparatively non-trivial to average latent state trajectories across multiple LDS models that are independently fit to data in overlapping time windows.
In section 3, we demonstrate more sophisticated examples where ${\bm{u}}_{t}^{k}$ is specified to track a continuously measured behavioral variable (e.g. heading direction or position of an animal) or follow a stepping or ramping function aligned to discrete task events (e.g. a sensory “go cue” or movement onset). Section 4 discusses further connections between CLDS models and existing state space models.
@@ 2.3 Inference
As mentioned earlier, the conditional distribution in eq. (2) has the advantage of describing a latent linear dynamical system (LDS), or linear Gaussian state-space model. As such, we can benefit from analytic tools like Kalman filtering to compute the filtering distributions $p({\mathbf{x}}_{t}\mid{\mathbf{y}}_{1:t},{\mathbf{F}},{\bm{u}}_{1:T})$ and marginal log-likelihood $p({\mathbf{y}}_{1:T}\mid{\mathbf{F}},{\bm{u}}_{1:T})$, and Kalman smoothing to compute the smoothing posterior $p({\mathbf{x}}_{1:T}\mid{\mathbf{y}}_{1:T},{\mathbf{F}},{\bm{u}}_{1:T})$. We focus on performing maximum-a-posteriori (MAP) inference for these parameters. In principle, it would be a straightforward extension to use variational inference or Markov Chain Monte Carlo to approximate the full posterior over these parameters.
@@
${\mathbf{y}}_{n}$ $={\mathbf{M}}({\bm{u}}_{n}){\mathbf{x}}_{n}+\bm{\epsilon}_{n},\quad\bm{\epsilon}_{n}\sim\mathcal{N}(0,\Sigma)$ (3a)
${\mathbf{M}}(\cdot)$ $\sim\mathcal{GP}^{D_{1}\times D_{2}}(0,k_{u}),$ (3b)
@@
Our parameterization in eq. (1c) implies that each entry is a dot product, ${\mathbf{M}}_{ij}(\cdot)={\bm{w}}^{(ij)\top}\bm{\phi}(\cdot)$, where $\bm{\phi}(\cdot)=(\phi_{1}(\cdot),\dots,\phi_{L}(\cdot))^{\top}\in\mathbb{R}^{L}$ is our vector of basis-functions evaluations. Therefore,
${\mathbf{M}}({\bm{u}})$ $={\mathbf{W}}^{\top}(\bm{\phi}({\bm{u}})\otimes{\bm{I}}_{D_{2}})$ (4)
${\mathbf{M}}({\bm{u}}){\bm{X}}$ $={\mathbf{W}}^{\top}(\bm{\phi}({\bm{u}})\otimes{\bm{X}}),$ (5)
for ${\bm{u}}\in\mathcal{U}$ and for any vector or matrix ${\bm{X}}\in\mathbb{R}^{D_{2},\dots}$ of appropriate dimension, with $\otimes$” the Kronecker product, and with our weights aggregated into the matrix
${\mathbf{W}}_{j+\ell,i}\coloneqq{\textnormal{w}}^{(ij)}_{\ell},\quad{\mathbf{W}}\in\mathbb{R}^{D_{2}L\times D_{1}}.$ (6)
With this, we can rewrite our regression problem as
@@
which is a linear regression problem with regularization from the prior $\log p({\mathbf{W}})=-\frac{1}{2}\left\|{\mathbf{W}}\right\|_{F}^{2}$ (up to an additive constant). We can analytically solve for the solution (derivations in Appendix §A.1.1), which yields that ${\mathbf{W}}_{\mathrm{MAP}}$ is the solution to the Sylvester equation
@@
Figure 2: Head direction synthetic experiment. (a) Schematic of latent dynamics and neural activity about $\theta\in[0,2\pi)$, the mouse HD, serving as conditions ${\bm{u}}=\theta$ in this task. (b) True nonlinear flow-field corresponding to the schematic in a, computed considering $p(\theta\mid{\mathbf{x}})=\delta_{\angle{\mathbf{x}}}(\theta)$. (c) Recovered composite dynamics by CLDS, see text for computation details. Grey scale indicates ${\mathbf{x}}_{t}$ occupancy. The model fixed points (colored) as a function of $\theta$ form a perfect ring, overlapping with the true fixed points. (d-e) Parameter recovery for the dynamics matrix ${\mathbf{A}}(\theta)$ (d) and the bias ${\mathbf{b}}(\theta)$ (e) as functions of head direction $\theta$. (f) Recovered eigenvalues of ${\mathbf{A}}(\theta)$ as a function of $\theta$, true in dashed. (g) Co-smoothing reconstruction from the test-set. The firing rate of one neuron is held-out (bottom) while the rest (top) is observed, and we reconstruct accurately the single-neuron firing rates for both the held-in (top) and held-out (bottom) neurons.
Expectation Maximization
We can leverage the above to perform MAP inference for $\{{\mathbf{A}},{\mathbf{b}},{\mathbf{C}},{\mathbf{d}},{\mathbf{m}}\}$ with the Expectation-Maximization (EM) algorithm (Dempster et al., 1977; Ghahramani & Hinton, 1996). In the E-step we obtain estimates of the moments of the latents with Kalman-smoothing, which then place us in a setting akin to eq. (3) with sufficient statistics as data and regressors. We can then perform closed-form M-steps with our updates in eq. (9). We provide in Appendix §A.1.2 an example of the associated derivations with these E- and M-steps for the joint update for ${\mathbf{A}}(\cdot)$ and ${\mathbf{b}}(\cdot)$. The resulting EM algorithm has several advantages: (1) all E- and M-steps are analytic, (2) the E-step provides us with exact (penalized) marginal log-likelihood calculations, and (3) the algorithm gives monotonic gradient ascent guarantees of the marginal log-likelihood (resp. log posterior) objective.
We initialize the EM algorithm at samples from our GP priors for the parameters. With the EM algorithm we also learn the covariance parameters $\{Q_{1},Q,R\}$. Finally, the hyper-parameters $\{L,\kappa,\sigma\}$ from the GP priors are determined through performance on held-out test sets from 80/20 trial splits on all experiments.
@@
The closed form EM updates are only applicable when the distribution of ${\mathbf{y}}_{t}$ conditioned on ${\mathbf{x}}_{t}$ and ${\bm{u}}_{t}$ is Gaussian. This assumption, stated in eq. (1), is common existing methods (e.g. Yu et al. 2009; Williams et al. 2018). However, alternative models are likely a better fit to many neural datasets. For spike count data, past work has utilized Poisson (e.g. Macke et al. 2011) and COM-Poisson (Stevenson, 2016) likelihoods.
Although we assume a Gaussian likelihood in our experiments in §3, we note that inference in CLDS models remains tractable whenever $p({\mathbf{y}}_{t}\mid{\mathbf{W}},{\mathbf{x}}_{t},{\bm{u}}_{t})$ is log concave. This condition satisfied by most likelihood models of interest (e.g. Poisson). Indeed, conditioned on ${\bm{u}}_{1:T}$, the log posterior density associated with ${\mathbf{x}}_{1:T}$ is equal to a sum of concave terms up to an additive constant (Paninski et al., 2010). Thus, we can use standard optimization routines to identify a MAP estimate of ${\mathbf{x}}_{1:T}$ efficiently with theoretical guarantees. This can then be used to implement an approximate EM algorithm (Macke et al., 2011).

### v2

@@ 1 Introduction
A central problem in neuroscience is to capture how neural dynamics are affected by external sensory stimuli, task variables, and behavioral covariates. To address this, a longstanding line of research has focused on characterizing neural dynamics through recurrent neural networks (RNNs) and their probabilistic counterparts, state-space models (SSMs; for reviews, see paninski2010new ; duncker2021dynamics ; Durstewitz2023 ).
Early work in this area utilized latent linear dynamical systems (LDS) with Gaussian observation noise. Although these assumptions are restrictive, they are beneficial in two respects. First, they simplify probabilistic inference by enabling Kalman smoothing and expectation maximization (EM)—two classical and highly effective methods GhahramaniHinton:EM . Second, they produce models that are mathematically tractable to analyze with well-established tools from linear systems theory (kailath1980linear, ). Indeed, many influential results in theoretical neuroscience have come from purely linear models seung1996brain ; goldman2009memory ; murphy2009balanced .
In reality, most neural circuits do not behave like time-invariant linear systems. Thus, more recent work from the machine learning community has cataloged a variety of nonlinear models for neural dynamics. Although these new models often predict held out neural data more accurately than LDS models, they are generally more difficult to fit and more difficult to understand. Thus, there has been a proliferation of competing models for neural data analysis, such as RNNs LFADS , transformers ye2021representation ; azabou2023a , and diffusion-based methods (kapoor2024latent, ), as well as competing inference methods, such as generalized teacher forcing Hess2023 , amortized variational inference LFADS , and sequential Monte Carlo pals2024inferring . Choosing among these strategies and scientifically interpreting the outcomes is challenging.
@@
Finally, CLDS models have connections to several existing methods (§4). For example, they can be viewed as a dynamical extension of a Wishart process model (Nejatbakhsh:2023, ) and an extension of Gaussian Process Factor Analysis (GPFA; Yu2009 ) with a learnable kernel and readout function that can vary across time and conditions. They are also similar to various forms of switching linear dynamical systems models (Petreska2011, ; nassar2018treestructured, ; hu2024modeling, ). The key difference is that the “switching” in CLDS models is governed by an observed covariate vector, ${\bm{u}}_{t}$, rather than by a discrete latent process. This makes inference in the CLDS model much more straightforward, albeit at the price of not being fully unsupervised.
@@ 2 Methods
We use ${\mathbf{f}}(\cdot)\sim\mathcal{GP}^{N}(\bm{m}(\cdot),k(\cdot,\cdot))$ with mean $\bm{m}:\mathcal{X}\to\mathbb{R}^{N}$ and kernel $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, to denote samples ${\mathbf{f}}:\mathcal{X}\to\mathbb{R}^{N}$ obtained from stacking independent Gaussian processes into an $N$-dimensional vector. Also, ${\bm{I}}_{D}$ is the $D\times D$ identity matrix.
@@ 2.1 Conditionally Linear Dynamical Systems
Consider an experiment with $N$ neurons recorded over $K$ trials of length $T$. Our dataset consists of the recorded neural firing-rate trajectories $\{{\mathbf{y}}_{1:T}^{k}\}_{k=1}^{K}$, with ${\mathbf{y}}_{t}\in\mathbb{R}^{N}$ at time-step $t\in\{1,\dots,T\}$, along with the corresponding experimental conditions $\{{\bm{u}}_{1:T}^{k}\}_{k=1}^{K}$, with ${\bm{u}}_{t}$ in the condition space $\mathcal{U}$. All modeling is done for single trials, and we drop superscripts $k$ where clear henceforth. By experimental conditions, we refer to available neural data covariates, either experimentally set or collected as measurements. These conditions, see Fig. 1a, can vary over time or remain constant (compare §3.4 vs. §3.5). They can also be a step function over time (e.g. animal moving vs. not moving), resulting in a switching-like mechanism between different dynamics.
We model responses ${\mathbf{y}}_{t}$ as emissions from a latent time-varying linear dynamical system in ${\mathbf{x}}_{t}\in\mathbb{R}^{D}$, with dynamics governed by the conditions ${\bm{u}}_{t}$. Specifically,
@@
over time steps $t\in\{1,\dots,T\}$ from initial condition ${\mathbf{x}}_{1}\sim\mathcal{N}\left({\mathbf{x}}_{1};{\mathbf{m}}({\bm{u}}_{1}),Q_{1}\right)$, and where $\bm{\epsilon}_{t}\sim\mathcal{N}(0,Q)$ and $\bm{\omega}_{t}\sim\mathcal{N}(0,R)$ are sources of noise, i.i.d. in time. We assume that the latent variables ${\mathbf{x}}_{t}$ follow smooth dynamics defined by time-varying linear matrices ${\mathbf{A}}({\bm{u}})\in\mathbb{R}^{D\times D}$ from initial mean ${\mathbf{m}}({\bm{u}}_{1})\in\mathbb{R}^{D}$, with bias terms ${\mathbf{b}}({\bm{u}})\in\mathbb{R}^{D}$, and emissions governed by ${\mathbf{C}}({\bm{u}})\in\mathbb{R}^{N\times D}$ and ${\mathbf{d}}({\bm{u}})\in\mathbb{R}^{N}$, all dependent on ${\bm{u}}\in\mathcal{U}$. See graphical depiction in Fig. 1b.
The system in (1a-1b) parameterizes a family of linear systems indexed by a continuous variable ${\bm{u}}_{t}$, which importantly is observed. Our approach can be viewed as the linearization in ${\mathbf{x}}_{t}$ of a fully nonlinear system in ${\mathbf{x}}_{t}$ and ${\bm{u}}_{t}$, under additive noise—we detail this relationship in Appendix §A.2. Most methods treat conditions as additive inputs, influencing the dynamics in (1a) via additive terms of the form ${\bm{B}}{\bm{u}}_{t}$ for a linear encoding matrix ${\bm{B}}\in\mathbb{R}^{D\times\left|\mathcal{U}\right|}$. In contrast, the mapping of experimental conditions onto linear dynamics, ${\bm{u}}\mapsto\{{\mathbf{A}}({\bm{u}}),{\mathbf{b}}({\bm{u}}),{\mathbf{C}}({\bm{u}}),{\mathbf{d}}({\bm{u}}),{\mathbf{m}}({\bm{u}})\}$, is allowed to be nonlinear and learnable. Specifically, we place an approximate Gaussian Process (GP) prior on each entry of the parameters through a finite expansion of basis functions, leveraging regular Fourier feature approximations (JMLR:v18:16-579, ). For any ${\mathbf{M}}\in\{{\mathbf{A}},{\mathbf{b}},{\mathbf{C}},{\mathbf{d}},{\mathbf{m}}\}$, we consider a prior
@@
over each $i,j$-th entry, truncated at $L\in\mathbb{N}$ basis functions. Each basis function $\phi_{\ell}:\mathcal{U}\to\mathbb{R}$ is fixed, and the randomness in the prior purely comes from the weights, ${\textnormal{w}}_{\ell}^{(ij)}$, which are drawn from a standard normal distribution. When constructed appropriately, the prior in equation (1c) converges to a nonparametric Gaussian process in the limit that $L\rightarrow\infty$, with kernel structure determined by the basis functions. Many choices are valid, and we elect to choose basis functions $\{\phi_{\ell}\}_{\ell=1}^{L}$ with the goal of expressivity; they are designed to approximate a GP prior of the form ${\mathbf{M}}_{ij}(\cdot)\sim\mathcal{GP}(0,k_{u})$ for the squared exponential kernel $k_{u}$ with variance $\sigma^{2}$ and length-scale $\kappa$ (NEURIPS2020_92bf5e62, , eq. 13).
We denote ${\mathbf{F}}=\{{\mathbf{A}},{\mathbf{b}},{\mathbf{C}},{\mathbf{d}},{\mathbf{m}}\}$ as the set of random functions, and analogously the parameter set ${\mathbf{F}}({\bm{u}})=\{{\mathbf{A}}({\bm{u}}),{\mathbf{b}}({\bm{u}}),{\mathbf{C}}({\bm{u}}),{\mathbf{d}}({\bm{u}}),{\mathbf{m}}({\bm{u}})\}$ for any condition ${\bm{u}}\in\mathcal{U}$. The model distribution
$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{1:T}\mid{\mathbf{A}},{\mathbf{b}},{\mathbf{C}},{\mathbf{m}},{\bm{u}}_{1:T})=p({\mathbf{y}}_{1:T},{\mathbf{x}}_{1:T}\mid{\mathbf{F}},{\bm{u}}_{1:T})$ (2)
describes a time-varying LDS, conditioned on a parameter sequence set at experimental conditions. Therefore, we refer to the model (1) as a Conditionally Linear Dynamical System (CLDS [^3] Our CLDS implementation is available at https://github.com/neurostatslab/clds.).
@@ 2.2 CLDS modeling choices
A second source of flexibility is the encoding of experimental covariates, $\bm{u}$}. Recall our notation from section 2.1, that ${\bm{u}}_{t}^{k}$ represents experimental covariates at time $t\in\{1,\dots,T\}$ and trial $k\in\{1,\dots,K\}$. A simple, and broadly applicable, modeling approach would be to set ${\bm{u}}_{t}^{k}=t$. This achieves a time-varying LDS model in which the GP prior encodes smoothness over time. This is similar in concept to fitting a linear model to data over a sliding time window costa2019adaptive ; galgali2023residual . However, the CLDS formulation of this idea is fully probabilistic, which has several advantages. For example, one can use a single pass of Kalman smoothing to infer the distribution over the latent state trajectory, ${\mathbf{x}}_{1:T}^{k}$, within each trial. It is comparatively non-trivial to average latent state trajectories across multiple LDS models that are independently fit to data in overlapping time windows.
In section 3, we demonstrate more sophisticated examples where ${\bm{u}}_{t}^{k}$ is specified to track a continuously measured behavioral variable (e.g. head direction or position of an animal) or follow a stepping or ramping function aligned to discrete task events (e.g. a sensory “go cue” or movement onset). Section 4 discusses further connections between CLDS models and existing state space models.
@@ 2.3 Inference
As mentioned earlier, the conditional distribution in eq. (2) has the advantage of describing a latent LDS, or linear Gaussian state-space model, given estimates of $\mathbf{F}$}. As such, we can benefit from analytic tools like Kalman filtering to compute the filtering distributions $p({\mathbf{x}}_{t}\mid{\mathbf{y}}_{1:t},{\mathbf{F}},{\bm{u}}_{1:t})$ and marginal log-likelihood $p({\mathbf{y}}_{1:T}\mid{\mathbf{F}},{\bm{u}}_{1:T})$, and Kalman smoothing for $p({\mathbf{x}}_{t}\mid{\mathbf{y}}_{1:T},{\mathbf{F}},{\bm{u}}_{1:T})$ and the latents posterior mode $\hat{\mathbf{x}}_{1:T}$. We focus on performing maximum-a-posteriori (MAP) inference for these parameters. In principle, it would be a straightforward extension to use variational inference or Markov Chain Monte Carlo to approximate the full posterior over these parameters.
@@
${\mathbf{y}}_{n}$ $={\mathbf{M}}({\bm{u}}_{n}){\mathbf{x}}_{n}+\bm{\epsilon}_{n},\qquad\bm{\epsilon}_{n}\sim\mathcal{N}(0,\Sigma),\quad{\mathbf{M}}(\cdot)\sim\mathcal{GP}^{D_{1}\times D_{2}}(0,k_{u}),$ (3)
@@
Our approximate basis representation in eq. (1c) implies that each entry is a dot product, ${\mathbf{M}}_{ij}(\cdot)={\bm{w}}^{(ij)\top}\bm{\phi}(\cdot)$, where $\bm{\phi}(\cdot)=(\phi_{1}(\cdot),\dots,\phi_{L}(\cdot))^{\top}\in\mathbb{R}^{L}$ is our vector of basis-functions evaluations. Therefore, we have
${\mathbf{M}}({\bm{u}}){\bm{X}}={\mathbf{W}}^{\top}(\bm{\phi}({\bm{u}})\otimes{\bm{X}}),$
for any ${\bm{u}}\in\mathcal{U}$ and vector or matrix ${\bm{X}}\in\mathbb{R}^{D_{2},\dots}$ of appropriate dimension, with $\otimes$” the Kronecker product, and with our weights aggregated into the matrix ${\mathbf{W}}_{D_{2}\ell+j,i}\coloneqq{\textnormal{w}}^{(ij)}_{\ell}$, ${\mathbf{W}}\in\mathbb{R}^{D_{2}L\times D_{1}}$. Derivations are provided in Appendix §A.1.1. With this, we can rewrite our regression problem as
@@
which is a linear regression problem with regularization from the prior $\log p({\mathbf{W}})=-\frac{1}{2}\left\|{\mathbf{W}}\right\|_{F}^{2}$ (up to an additive constant) from (1c). We can analytically solve for the solution (derivations in Appendix §A.1.2), which yields that ${\mathbf{W}}_{\mathrm{MAP}}$ is the solution to the Sylvester equation
@@
Figure 2: Head direction synthetic experiment. (a) Schematic of latent dynamics and neural activity about $\theta\in[0,2\pi)$, the mouse HD, serving as conditions ${\bm{u}}=\theta$ in this task. (b) True nonlinear flow-field corresponding to the schematic in a, computed considering $p(\theta\mid{\mathbf{x}})=\delta_{\angle{\mathbf{x}}}(\theta)$. (c) Recovered composite dynamics by CLDS (see text). Grey scale indicates posterior $\hat{\mathbf{x}}_{t}$ occupancy, and thus our ability to estimate the dynamics. The model fixed points (colored) as a function of $\theta$ form a perfect ring, overlapping with the true fixed points. (d-e) Parameter recovery for the dynamics matrix ${\mathbf{A}}(\theta)$ (d) and the bias ${\mathbf{b}}(\theta)$ (e) as functions of head direction $\theta$. (f) Recovered eigenvalues of ${\mathbf{A}}(\theta)$ as a function of $\theta$, true in dashed. (g) Co-smoothing reconstruction from the test-set. The firing rate of one neuron is held-out (bottom) while the rest (top) is observed, and we reconstruct accurately the single-neuron firing rates for both the held-in (top) and held-out (bottom) neurons.
Inference with Expectation Maximization
We can leverage the above to perform MAP inference for ${\mathbf{F}}=\{{\mathbf{A}},{\mathbf{b}},{\mathbf{C}},{\mathbf{d}},{\mathbf{m}}\}$ with the Expectation-Maximization (EM) algorithm (Dempster1977, ; GhahramaniHinton:EM, ). In the E-step we obtain estimates of the moments of the latents with Kalman-smoothing, which then place us in a setting akin to eq. (3) with sufficient statistics as data and regressors. We can then perform closed-form M-steps with our updates in eq. (6). We provide in Appendix §A.1.3 an example of the associated derivations with these E- and M-steps for the joint update for ${\mathbf{A}}(\cdot)$ and ${\mathbf{b}}(\cdot)$. The resulting EM algorithm has several advantages: (1) all E- and M-steps are analytic, (2) the E-step provides us with exact (penalized) marginal log-likelihood calculations, and (3) the algorithm gives monotonic gradient ascent guarantees of the marginal log-likelihood (resp. log posterior) objective.
We initialize the EM algorithm at samples from our GP priors for $\mathbf{F}$}. With the EM algorithm we also learn the covariance parameters $\{Q_{1},Q,R\}$. The hyper-parameters $\{L,\kappa,\sigma\}$ for the GP priors and the dimensionality of the latents $D$ are determined through performance on held-out test sets from 80/20 trial splits on all experiments unless specified.
@@
The closed form EM updates are only applicable when the observation distribution of ${\mathbf{y}}_{t}$ conditioned on ${\mathbf{x}}_{t}$ and ${\bm{u}}_{t}$ (1b) is Gaussian. This assumption is common to existing methods (e.g. Yu2009 ; williams2018unsupervised ), though to model spike count data directly, most work utilize Poisson (e.g. macke2011empirical ) and COM-Poisson (stevenson2016flexible, ) likelihoods. Such emission likelihoods of interest have log concave distributions, for which inference in CLDS models would remain tractable. Indeed, conditioned on ${\bm{u}}_{1:T}$, the log posterior density associated with ${\mathbf{x}}_{1:T}$ is equal to a sum of concave terms up to an additive constant (paninski2010new, ). Thus, we can use standard optimization routines to identify a MAP estimate of ${\mathbf{x}}_{1:T}$ with theoretical guarantees, and leverage it for approximate EM (macke2011empirical, ).